# Artificial Precision Timing Array: bridging the decihertz gravitational-wave sensitivity gap with clock satellites

### v1

@@ Abstract
Gravitational-wave astronomy has developed enormously over the last decade with the first detections across different frequency bands, but has yet to access 0.1100.1100.1–100.1 – 10 Hz gravitational waves. Gravitational waves in this band are emitted by some of the most enigmatic sources, including intermediate-mass binary black hole mergers, early inspiralling compact binaries, and possibly cosmic inflation. To tap this exciting band, we propose the construction of a detector based on pulsar timing principles, the Artificial Precision Timing Array (APTA). We envision APTA as a solar system array of artificial “pulsars”—precision-clock-carrying satellites that emit pulsing electromagnetic signals towards Earth or other centrum. In this fundamental study, we estimate the clock precision needed for APTA to successfully detect gravitational waves. Our results suggest that a clock relative uncertainty of $10^{-17}$, which is currently attainable, would be sufficient for APTA to surpass LISA’s sensitivity in the decihertz band and observe $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ black hole mergers. Future atomic clock technology realistically expected in the next decade would enable the detection of an increasingly diverse set of astrophysical sources, including stellar-mass compact binaries that merge in the LIGO-Virgo-KAGRA band, extreme-mass-ratio inspirals, and Type Ia supernovae. This work opens up a new area of research into designing and constructing artificial gravitational-wave detectors relying on the successful principles of pulsar timing.
@@ I Introduction
Since the first successful detection of gravitational waves (GWs) by LIGO in 2015, more than 100 compact-binary mergers have been seen by ground-based detectors Aasi et al. (2015); Abbott et al. (2016, 2021); Accadia et al. (2012). Moreover, the global network of pulsar timing arrays (PTAs) has provided strong evidence of the stochastic GW background Agazie et al. (2023a); Antoniadis et al. (2023); Reardon et al. (2023); Zic et al. (2023); Xu et al. (2023). With the fourth LIGO-Virgo-KAGRA observing run in progress and the continued PTA observations Manchester et al. (2013); Desvignes et al. (2016); Antoniadis et al. (2022); Agazie et al. (2023a), the era of gravitational-wave astronomy is surely underway.
@@
Current techniques access a limited window of the GW frequency spectrum. The ongoing PTA observation campaign is sensitive in the $10^{-9}–10^{-6}$ Hz band Maiorano et al. (2021), with its upper-frequency limit set by its fortnightly access to radio telescopes. Meanwhile, LIGO-Virgo-KAGRA interferometers are sensitive in the $\mathrm{10\ Hz–10\ kHz}$ band. In the future, space-based interferometers such as Laser Interferometer Space Antenna (LISA) and TianQin may be sensitive to 0.11000.11000.1–1000.1 – 100 mHz GWs, while next-generation Earth-based missions such as Cosmic Explorer and Einstein Telescope will push the lower limit of ground-based sensitivity to a few Hz Amaro-Seoane et al. (2017); Huerta and Gair (2010). Seismic and gravity gradient noises, however, severely limit ground-based detector sensitivity below $\sim$ 1 Hz Abbott et al. (2009); Punturo et al. (2010); Abbott et al. (2016). Consequently, the 0.110.110.1–10.1 – 1 Hz remains inaccessible, although space-based interferometer concepts, such as the DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) Kawamura et al. (2006) and the Big Bang Observer (BBO) Phinney et al. (2004); Harry et al. (2006), have been proposed.
@@
Observational access to the 0.1–10 Hz GW band is imperative for the progress of many areas of astrophysics and cosmology. For example, $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ binary black holes have yet to be unambiguously observed. The observation of these sources can advance our understanding of supermassive black hole and galaxy formation, stellar clusters, and active galactic nuclei (AGN) Miller and Colbert (2004); Milosavljević et al. (2009); McKernan et al. (2012); Ferrara et al. (2014); Yang et al. (2019). While LISA may detect the early inspirals of intermediate-mass binaries Jani et al. (2020), a 0.1–10 Hz detector accesses the wealth of additional information only available during a merger. Other promising phenomena that may feature in this sensitivity gap include primordial GWs from inflation Corbin and Cornish (2006); Smith et al. (2006); Kuroyanagi et al. (2009); Guzzetti et al. (2016), the evolution of stellar-mass binaries and merging triple black hole systems Arca Sedda et al. (2021); O’Leary et al. (2009); Samsing (2018); Gondán et al. (2018); Antonini et al. (2017); Samsing and D’Orazio (2018); Codazzo et al. (2023), supernovae Ott (2009); Seitenzahl et al. (2015), and parabolic encounters Kocsis et al. (2006); Codazzo et al. (2023), among others further discussed in Section II.
@@
To reach this unexplored GW band, we propose a new detector concept: the Artificial Precision Timing Array (APTA), a network of satellites distributed around the Earth carrying high-precision clocks to act as artificial “pulsars”. Having satellite signals aimed at a small station (or stations) dedicated to their reception, APTA can avoid the cadence constraint of regular PTAs and achieve high sensitivity at much higher frequencies. Then, if the satellite clocks are timed at least every $\sim$ 0.1 s, APTA can reach high sensitivity to 0.1–10 Hz GWs for observations lasting 10 s or more.
@@
The idea of using high-precision clocks in space to detect GWs has been proposed and applied before. The global positioning system (GPS) was used to constrain the strain amplitude of the GW background in the 0.01–1 Hz band to $h_{c}<4.8\times 10^{-12}(1~{}{\rm Hz}/f)$ Aoyama et al. (2014), in an application of the Doppler tracking method Estabrook and Wahlquist (1975); Armstrong (2006); Vutha (2015). Another study proposed the use of two satellites in heliocentric orbits equipped with optical lattice atomic clocks to detect GWs, using a technique similar to Doppler tracking Kolkowitz et al. (2016). Another proposed detector, the Space Atomic Gravity Explorer (SAGE), consists of two satellites carrying ultracold atomic strontium systems to either perform atom interferometry with the strontium or Doppler tracking utilizing Sr optical atomic clocks Tino et al. (2019). Nevertheless, the idea of equipping artificial “pulsars” in the solar system with high-precision clocks to detect GWs with pulsar-timing-type methods is unprecedented.
@@
In this paper, we describe the groundwork for a solar system APTA targeting the 0.1–10 Hz GW frequency band. In Section II, we list and discuss several interesting astrophysical targets observable in this band. In Section III, we summarize the basics of PTA sensitivity Moore et al. (2015a). In Section IV, we present the concept of APTA in detail, including its sensitivity for different configurations and potential performance. In Section V, we briefly list and discuss error-introducing factors to be incorporated into a more realistic future model of APTA. We conclude in Section VI.
II Gravitational Wave Astrophysics with APTA
@@
Inspiraling stellar-mass compact binaries also emit GWs in this band before merging in the ground-accessible band. Low-frequency GW observations of these sources directly probe their formation and evolution Nishizawa et al. (2016, 2017); Arca Sedda et al. (2021); O’Leary et al. (2009); Samsing (2018); Gondán et al. (2018); Antonini et al. (2017). Long-duration tracking may probe BBH eccentricity (Arca Sedda et al., 2020; Liu et al., 2022) as well as improve effective spin and localization measurements. Similarly the formation of merging triple black hole systems may also be probed in this band Samsing and D’Orazio (2018); Codazzo et al. (2023). Excitingly, early detection of an inspiraling compact binary could trigger multimessenger precursor counterpart searches Tsang et al. (2012); Sullivan et al. (2023) as well as give advanced warning of short gamma-ray bursts (Narayan et al., 1992; Abbott et al., 2017a, b) and kilonovae (see Ref. Gottlieb et al. (2023) and references therein).
The characteristic GW strain of an astrophysical binary well before merger may be estimated as Jaffe and Backer (2003); Finn and Thorne (2000a)
$h_{c}\approx 4.1\times 10^{-20}\left(\frac{M}{10^{3}\text{ M}{{}_{\odot}}}\right)^{\frac{5}{6}}\left(\frac{f_{GW}}{1\text{ Hz}}\right)^{-\frac{1}{6}}\left(\frac{D}{1\text{ Gpc}}\right)^{-1}\frac{q^{\frac{1}{2}}}{(1+q)},$ (1)
where $M$ is the total mass, $f_{GW}$ is the gravitational wave frequency, $D$ is the luminosity distance to the source, and $q\leq 1$ is the mass ratio. The merger frequency for a BBH may be estimated as
@@
$f_{GW}^{merger}\approx\frac{1}{\pi}\frac{c^{3}}{2^{\frac{3}{2}}GM}\approx 1.3\times 10^{4}\text{ Hz}\left(\frac{M}{1\text{ M}_{\odot}}\right)^{-1}.$ (2)
@@
Heavy stellar-mass BBHs (total mass $\sim 100$ $\mathrm{M}_{\odot}$) at distances of $\sim 1$ Gpc will have strains in the range $h\sim 10^{-21}–10^{-20}$ in dHz band during early inspiral. The late inspirals and mergers of $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ BBHs should have even larger dHz GW strains. For instance, an intermediate-mass binary black hole merger with $D\sim 1$ Gpc, and total mass $\sim 1\times 10^{4}$ ${}_{\odot}$, which will merge at $f_{GW}\approx 1.3$ Hz, will have strains exceeding $h\gtrsim 10^{-19}$. Since the GW frequency is half the orbital frequency, the time over which massive merging BBHs will emit 0.1-10 Hz GWs may be estimated as Peters (1964)
$T\approx 70\text{ s}\left(\frac{M}{10^{4}\text{ M}}_{\odot}\right)^{-\frac{5}{3}}\left(\frac{f_{GW}}{0.01\text{ Hz}}\right)^{-\frac{8}{3}}\frac{(1+q)^{2}}{q}.$ (3)
@@
$10^{3}–10^{4}$ ${}_{\odot}$ will spend $100-10^{4}$ s in the APTA band before merging, while stellar-mass BBHs can remain for days to even years depending on the mass, allowing for long-term signal accumulation from these sources.
@@
More unique astrophysical GW sources may also manifest in this frequency regime. Low frequency sources such as Type Ia Seitenzahl et al. (2015) and core-collapse supernovae Ott (2009), parabolic encounters Kocsis et al. (2006); Codazzo et al. (2023), white dwarf-neutron star mergers (Rebassa-Mansergas et al., 2019; Korol et al., 2022), and even some extreme-mass-ratio inspirals Finn and Thorne (2000b); Hughes (2000); Babak et al. (2017) should emit GWs in this band. With longer observations, APTA can tap lower frequencies. As such, APTA could be able to observe verification binaries to be used in LISA calibration Kupfer et al. (2018); Marsh (2011); Korol et al. (2017); Breivik et al. (2018); Tauris (2018), as well as probe the major astrophysics goals of LISA Amaro-Seoane et al. (2023).
@@
Observations in the 0.1–10 Hz frequency band could also help constrain models of cosmic inflation through the detection of primordial gravitational waves Corbin and Cornish (2006); Smith et al. (2006); Kuroyanagi et al. (2009); Guzzetti et al. (2016). It is advantageous to explore this frequency range as opposed to LISA’s 0.11000.11000.1–1000.1 – 100 mHz Amaro-Seoane et al. (2017), since, in the latter, inflationary GWs may be overshadowed by binary white dwarf GWs, which disappear entirely above 0.25 Hz Ungarelli and Vecchio (2001); Farmer and Phinney (2003); Cutler and Holz (2009). Furthermore, heavy ($\lesssim 1$ $\mathrm{M}_{\odot}$) primordial-black-hole binaries, which could account for a small fraction of the dark matter in the form of massive compact halo objects (MACHOs), could be detected in the APTA band during their early inspiral Nakamura et al. (1997); Raidal et al. (2019). Another possibility of cosmological interest in the APTA band could be the direct measurement of the acceleration of the expansion of the universe and, therefore, the curvature parameter $\Omega_{k0}$ Nakamura and Chiba (1999); Seto et al. (2001).
III Pulsar Timing Array Sensitivity

### v2

@@ Abstract
Gravitational-wave astronomy has developed enormously over the last decade with the first detections and continuous development across broad frequency bands. However, the decihertz range has largely been left out of this development. Gravitational waves in this band are emitted by some of the most enigmatic sources, including intermediate-mass binary black hole mergers, early inspiraling compact binaries—whose late evolution and merger are seen by Earth-based detectors—, and possibly primordial gravitational waves. To tap this exciting band, we propose the construction of a detector based on precision timing principles, the Artificial Precision Timing Array (APTA). We envision APTA as a solar system array of artificial “pulsars”—precision-time-reference-carrying satellites that emit periodic electromagnetic signals towards Earth or other satellite constellation centrum. In this fundamental study, we estimate the clock precision needed for APTA to be able to detect gravitational waves. Our results suggest that 6 satellites and a clock relative uncertainty of $10^{-18}$ at 1 s of averaging, which is currently attainable with atomic clocks, would be sufficient for APTA to reach pristine sensitivity in the decihertz band and be sensitive to $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ black hole mergers and the early inspiral of heavy LIGO-Virgo-KAGRA sources. Future time reference, oscillator, and clock technologies realistically expected in the next decade(s) would enable the detection of an increasingly diverse set of astrophysical sources and allow APTA to reach a better sensitivity than other detector concepts proposed for the decihertz band. This work opens up a new area of research into designing and constructing gravitational-wave detectors relying on principles used successfully in pulsar timing.
@@ I Introduction
Since the first detection of gravitational waves (GWs) by LIGO in 2015 Aasi et al. (2015); Abbott et al. (2016, 2021a); Abac et al. (2025), more than 300 compact-binary mergers have been seen by ground-based detectors. Moreover, the global network of pulsar timing arrays (PTAs) has provided strong evidence for a stochastic GW background Accadia et al. (2012); Akutsu et al. (2019); Agazie et al. (2023a); Antoniadis et al. (2023); Reardon et al. (2023); Zic et al. (2023); Xu et al. (2023). With the fourth LIGO-Virgo-KAGRA observing run in progress and the continued PTA campaign Manchester et al. (2013); Accadia et al. (2012); Antoniadis et al. (2022); Agazie et al. (2023a), more and more observational data and, thus, scientific insight can be expected.
While Earth-based detectors utilize interferometry to detect GWs, PTAs, including NANOGrav (Agazie et al., 2023a), the European Pulsar Timing Array Desvignes et al. (2016); Antoniadis et al. (2023), the Parkes Pulsar Timing Array Manchester et al. (2013); Reardon et al. (2023); Zic et al. (2023), the Chinese Pulsar Timing Array Xu et al. (2023), and the MeerKAT Pulsar Timing Array Bailes et al. (2020), rely on the pulsar timing technique. Pulsars, rapidly rotating magnetized neutron stars, emit regular radio pulses with periods ranging from $10^{-3}$ to 10 s Condon and Ransom (2016). The change in their spin periods $\dot{P}$ is $\dot{P}\lesssim 10^{-19}$ ss-1 Lorimer (2008). With these very precisely measured and stable periods over long timescales, millisecond pulsars make for some of the powerful natural clocks. A GW passing between a pulsar and the Earth alters the travel time of the radio pulse from the star to the telescope. Subtracting a model for the times of arrival (ToAs) from the observed data yields a set of timing residuals, deviations from the expected ToAs, which may come from GWs or noise processes.
@@
Current techniques access a limited window of the GW frequency spectrum. The ongoing PTA observation campaign is sensitive in the $10^{-9}–10^{-6}$ Hz band Maiorano et al. (2021). Usually, PTAs have their upper-frequency limit set by their fortnightly access to radio telescopes. Even when this practical concern is overcome, there is a fundamental frequency upper limit of order 1 mHz due to the folding time required to process pulsar observations Dolch et al. (2014, 2016). Meanwhile, LIGO-Virgo-KAGRA interferometers are sensitive in the $\mathrm{10~Hz–10~kHz}$ band. In the future, space-based interferometers such as the Laser Interferometer Space Antenna (LISA) and TianQin may be sensitive to 0.11000.1–100 mHz GWs, while next-generation Earth-based missions such as Cosmic Explorer (CE) and Einstein Telescope (ET) will push the lower limit of ground-based sensitivity to a few Hz Luo et al. (2016); Amaro-Seoane et al. (2017); Reitze et al. (2019); Huerta and Gair (2010). Still, seismic and gravity gradient noises severely limit ground-based detector sensitivity below $\sim$ 1 Hz Abbott et al. (2009); Punturo et al. (2010); Abbott et al. (2016). Consequently, the 0.110.1–1 Hz region remains inaccessible, although nice concepts, such as the Advanced Laser Interferometer Antenna (ALIA) Bender (2004); Bender and Begelman (2005), the Big Bang Observer (BBO) Phinney et al. (2004); Harry et al. (2006), the DECi-hertz Interferometer Gravitational wave Observatory (DECIGO) Kawamura et al. (2006), the Gravitational-Wave Lunar Observatory for Cosmology (GLOC) Jani and Loeb (2020), and the Lunar Gravitational-wave Antenna (LGWA) Ajith et al. (2024) have been proposed.
@@
Observational access to the 0.1–10 Hz GW band is imperative for the progress of many areas of astrophysics and cosmology. For example, $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ binary black holes (BBHs) have yet to be unambiguously observed. The observation of these sources can advance our understanding of supermassive black hole and galaxy formation, stellar clusters, and active galactic nuclei (AGN) Miller and Colbert (2004); Milosavljević et al. (2009); McKernan et al. (2012); Ferrara et al. (2014); Yang et al. (2019). While LISA may detect the early inspirals of $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ binaries Jani et al. (2020), a 0.1–10 Hz detector accesses the wealth of additional information only available during their mergers. Other promising phenomena that may feature in this band include primordial GWs Corbin and Cornish (2006); Smith et al. (2006); Kuroyanagi et al. (2009); Guzzetti et al. (2016), the evolution of stellar-mass binaries and triple black hole systems Arca Sedda et al. (2021); O’Leary et al. (2009); Samsing (2018); Gondán et al. (2018); Antonini et al. (2017); Samsing and D’Orazio (2018); Codazzo et al. (2023), supernovae Ott (2009); Seitenzahl et al. (2015), and parabolic encounters Kocsis et al. (2006); Codazzo et al. (2023), among others further discussed in Section II.
@@
To reach this unexplored GW band, we propose a new detector concept: the Artificial Precision Timing Array (APTA), a network of satellites distributed around the Earth or another centrum carrying high-precision clocks or comparably accurate time references to act as artificial “pulsars”. Having APTA clock satellite signals be clearly modulated and aimed at a small station (or stations) dedicated to their reception, APTA can avoid the cadence constraint of regular PTAs and achieve high sensitivity at much higher frequencies. If the satellite signals are pulsed at more than 10 Hz, APTA can reach high sensitivity to 0.1–10 Hz (the APTA band) GWs for observations lasting more than 10 s.
@@
The idea of using high-precision clocks in space for the empirical study of GWs has been proposed and applied before. The global positioning system (GPS) was used to constrain the strain amplitude of the GW background in the 0.01–1 Hz band Aoyama et al. (2014), in an application of the Doppler tracking method Estabrook and Wahlquist (1975); Armstrong (2006); Vutha (2015). Another study proposed the use of two satellites in heliocentric orbits equipped with optical lattice atomic clocks to detect GWs, using a technique similar to Doppler tracking Kolkowitz et al. (2016). Another proposed detector, the Space Atomic Gravity Explorer (SAGE), consists of two satellites carrying ultracold atomic strontium systems to either perform atom interferometry with the strontium or Doppler tracking utilizing Sr optical atomic clocks Tino et al. (2019). Nevertheless, the idea of performing pulsar-timing-type GW detection with artificial satellites equipped with high-precision clocks is unprecedented.
@@
In this paper, we describe the groundwork for APTA targeting the 0.1–10 Hz GW frequency band. In Section II, we list and discuss several interesting astrophysical targets observable in this band. In Section III, we derive a characteristic strain sensitivity curve for APTA. In Section IV, we present the concept of APTA in detail, discussing parameter choices for the detector and showing its sensitivity curves alongside other detectors and sources. In Section V, we briefly list and discuss error-introducing factors to be incorporated into a more realistic future model of APTA. We conclude in Section VI.
II Gravitational-Wave Astrophysics with APTA
@@
Inspiraling stellar-mass compact binaries also emit GWs in this band before merging in the ground-accessible band. Low-frequency GW observations of these sources directly probe their formation and evolution Nishizawa et al. (2016, 2017); Arca Sedda et al. (2021); O’Leary et al. (2009); Samsing (2018); Gondán et al. (2018); Antonini et al. (2017). Long-duration tracking may probe BBH eccentricity (Arca Sedda et al., 2020; Liu et al., 2022) as well as improve effective spin and localization measurements. Similarly, the formation of merging triple black hole systems may also be probed in this band Samsing and D’Orazio (2018); Codazzo et al. (2023). Excitingly, early detection of an inspiraling compact binary could trigger a wide range of observations in multimessenger astronomy, such as counterpart and precursor searches Tsang et al. (2012); Sullivan et al. (2023), as well as give advance warning for gravitational-wave detectors in higher frequency bands, gamma-ray-burst observatories (Narayan et al., 1992; Abbott et al., 2017a, b), and telescopes looking for kilonova or other multimessenger signatures (see Ref. Gottlieb et al. (2023) and references therein).
The characteristic GW strain of an astrophysical binary at redshift $z$ may be estimated as Sesana et al. (2005)
$h_{c}\approx 5.3\times 10^{-20}\left(\frac{f_{\rm gw}}{0.1\text{~Hz}}\right)^{-\frac{1}{6}}\left(\frac{M}{10^{3}\text{ M}{{}_{\odot}}}\right)^{\frac{5}{6}}\frac{q^{\frac{1}{2}}}{(1+q)}\\ \left(\frac{(1+z)}{1.2}\right)^{\frac{5}{6}}\left(\frac{D}{1\text{~Gpc}}\right)^{-1},$ (1)
where $M$ is the source-frame total mass, $f_{\rm gw}$ is the observed GW frequency, $q\leq 1$ is the mass ratio, and $D$ is the luminosity distance to the source, which directly determines $z$ given a set of cosmological parameters—which we take from Planck 2018 for this study Aghanim et al. (2020). The merger frequency for a BBH may be estimated with a Newtonian approximation as
@@
$f_{\rm gw}^{\rm merger}\approx\frac{1}{\pi}\frac{c^{3}}{2^{\frac{3}{2}}GM}\approx 2.3\times 10^{4}\text{~Hz}\left(\frac{M}{1\text{ M}_{\odot}}\right)^{-1}.$ (2)
@@
Heavy stellar-mass BBHs (total mass $\sim 100$ $\mathrm{M}_{\odot}$) at distances of $\sim 1$ Gpc will have strains of order $h_{c}\sim 10^{-21}$ in the dHz band during their early inspiral. The late inspirals and mergers of $10^{3}–10^{4}$ $\mathrm{M}_{\odot}$ BBHs should have even larger dHz GW strains. For instance, an intermediate-mass binary black hole merger with $D\sim 1$ Gpc, and total mass $\sim 1\times 10^{4}$ M, which will merge at $f_{\rm gw}\approx 2.3$ Hz, will have strains of $h_{c}\sim 10^{-19}$.
Since the GW frequency is twice the orbital frequency, the time in the detector frame over which massive merging BBHs will emit 0.1-10 Hz GWs may be estimated as Peters (1964)
$T\approx 70\text{~s }\left(\frac{1+z}{1.2}\right)\left(\frac{M}{10^{4}\text{ M}}_{\odot}\right)^{-\frac{5}{3}}\left(\frac{f_{\rm gw}}{0.1\text{~Hz}}\right)^{-\frac{8}{3}}\frac{(1+q)^{2}}{q}.$ (3)
@@
Therefore, $10^{3}–10^{4}$ M binaries will spend $10^{2}–10^{4}$ s in the APTA band before merging, while stellar-mass BBHs can remain for days to even years depending on the mass, allowing for long-term signal accumulation from these sources.
@@
More unique astrophysical GW sources may also manifest in this frequency regime. Low-frequency sources such as Type Ia Seitenzahl et al. (2015) and core-collapse supernovae Ott (2009), parabolic encounters Kocsis et al. (2006); Codazzo et al. (2023), white dwarf-neutron star mergers (Rebassa-Mansergas et al., 2019; Korol et al., 2022; Kang et al., 2024), and even some extreme-mass-ratio inspirals Finn and Thorne (2000); Hughes (2000); Babak et al. (2017) should emit GWs in this band. With longer observations, APTA can tap lower frequencies. As such, APTA could be able to observe verification binaries to be used in LISA calibration Kupfer et al. (2018); Marsh (2011); Korol et al. (2017); Breivik et al. (2018); Tauris (2018), as well as probe the major astrophysics goals of LISA Amaro-Seoane et al. (2023).
@@
Observations in the 0.1100.1–10 Hz frequency band could also help constrain models of early-universe physics through the detection of primordial gravitational waves Corbin and Cornish (2006); Smith et al. (2006); Kuroyanagi et al. (2009); Guzzetti et al. (2016). It is advantageous to explore this frequency range as opposed to LISA’s 0.11000.1–100 mHz Amaro-Seoane et al. (2017) since, in the latter, primordial GWs may be overshadowed by binary white dwarf GWs, which disappear entirely above 0.25 Hz Ungarelli and Vecchio (2001); Farmer and Phinney (2003); Cutler and Holz (2009). Furthermore, heavy ($\lesssim 1$ $\mathrm{M}_{\odot}$) primordial-black-hole (PBH) binaries, which could account for a small fraction of the dark matter in the form of massive compact halo objects (MACHOs), could be detected in the APTA band during their early inspiral Nakamura et al. (1997); Raidal et al. (2019). Another possibility of cosmological interest in the APTA band could be the direct measurement of the acceleration of the expansion of the universe and, therefore, the curvature parameter $\Omega_{k0}$ Nakamura and Chiba (1999); Seto et al. (2001).
III APTA Sensitivity
III.1 The Physics of Pulsar-Timing-Based GW Detection
In this subsection, we provide a theoretical overview of the physical mechanism underlying how PTAs, and hence APTA, work. Our calculation follows closely the one in Ref. Maggiore (2018) and we initially work in $c=1$ units.
Gravitational waves are linear perturbations to a flat background spacetime. In transverse-traceless (TT) gauge, suitable to study GW propagation in vacuum, the spacetime line element $ds$ is defined by
$ds^{2}=-dt^{2}+\left[\delta_{ij}+h_{ij}^{TT}(t,\vec{x})\right]dx^{i}dx^{j}.$ (4)
where $t$ is the time coordinate, $x^{i}$, $i\in\{1,2,3\}$, are spatial coordinates, $\delta_{ij}$ is the Kronecker delta and $h_{ij}^{TT}$ is the TT gravitational-wave tensor.
We consider a coordinate system with its origin at APTA’s signal reception station, in which one of its satellites is at position $\vec{r_{s}}=d_{s}\mathbf{\hat{x}}$ on the x-axis (the basic architecture of APTA is discussed in Section IV). Following the null (i.e. $ds^{2}=0$) trajectory $\vec{x}(t)$ of a photon emitted at time $t_{e}$ from the satellite to the centrum,
$dx^{2}=\frac{dt^{2}}{1+h^{TT}_{xx}(t,\vec{x}(t))}.$ (5)
The photon is traveling from positive $x$ to the origin, so we get the solution moving in the x-x direction:
$dx=\frac{-dt}{\sqrt{1+h^{TT}_{xx}(t,\vec{x}(t))}}\approx-\left[1-\frac{1}{2}h^{TT}_{xx}(t,\vec{x}(t))\right]dt$ (6)
to first order in the perturbation $h_{xx}^{TT}$.
When dealing with a single pulsar or APTA satellite, one can always choose the coordinate system so that the light source sits along one of the axes. However, to proceed to a result for a light source on an arbitrary direction $\mathbf{\hat{p}}$, which will be useful when we deal with multiple beacons, we replace $h_{xx}^{TT}$ with $\hat{p}^{i}\hat{p}^{j}h^{TT}_{ij}$.
Next, we can integrate from $t_{e}$ to the observation time $t_{o}$ so that